OpenStudy (anonymous):

Which measurement of length is the smallest? A. 9*10^-6 m B. 7*10^-5 m C. 7 * 10^1 D. 9*10^2

OpenStudy (somy):

what do u think?

OpenStudy (anonymous):

I have no idea.

OpenStudy (somy):

okay which grade are you in? and are you allowed to use calculator?

OpenStudy (anonymous):

Yes. And 11th.

OpenStudy (somy):

okay then write the numbers in calculator and u'll see which one is the smallest

OpenStudy (jfraser):

since all the measurements are in powers of 10, just find the one with the smallest exponent

OpenStudy (somy):

A. 9*10^-6 m = means has 6 zeros before 9 meaning 0.000009 m B. 7*10^-5 m = means has 5 zeros before 6 = 0.00007 m C. 7 * 10^1 m = here it means 1 zero after 7 = 70 m D. 9*10^2 m = and here its 2 zeros after 9 = 900 m

OpenStudy (somy):

i hope u get it?

OpenStudy (anonymous):

So would it be C?

OpenStudy (somy):

why do u think its C?

OpenStudy (somy):

take a look at ones that have decimal places, don't you think they are smaller?

OpenStudy (somy):

they are not even close to 1 right?

OpenStudy (anonymous):

Can you maybe help me with this question next? Scientists use scientific notation to represent very small or very large numbers because this notation increases the accuracy of measured quantities. convenience in using the numbers. number of significant figures. precision of measurements.

OpenStudy (somy):

maybe convenience in using the numbers.? @JFraser ?

OpenStudy (jfraser):

if you had to write 6.022*10^23 in "regular" notation, wouldn't that be terribly \(inconvenient\)?

OpenStudy (anonymous):

Ok thank you you!

OpenStudy (anonymous):

my calculator is messing up. I also need help with this. David writes the four numbers 6.91 10–6, –7.419 106, –9.71 10–5, and 6.42 106 on the board. In which option are these numbers ordered from least to greatest?

OpenStudy (somy):

can u write those numbers better?

OpenStudy (anonymous):

David writes the four numbers 1. 6.9*10^–6, 2. –7.419 * 10^6, 3. –9.71 8 10^ –5, and 4. 6.42 8 10^6 on the board. In which option are these numbers ordered from least to greatest?